1.    Given the utility function of the form: U (x, y) = 4x2 + 3xy + 6y2: maximize utility subject to the budget constraint: x + y = 56. Then find the utility maximizing level of output x and y?


1
Expert's answer
2021-09-23T09:14:30-0400

"U(x,y) = 4x^2 +3xy+6y^2"

Budget line

"56 = x+y \\

P(x)=1 \\

P(y) = 1"

At optimum : "\frac{MU(x)}{P(x)}= \frac{MU(y)}{P(y)}"

"MU(x) = \frac{ΔU}{Δx} = 8x+3y \\

MU(y) = \frac{ΔU}{Δy} = 3x+12y \\

\frac{8x+3y}{1} = \frac{3x+12y}{1} \\

8x+3y=3x+12y \\

5x=9y \\

x= \frac{9}{5}y = 1.8 y"

Putting x=1.8y in budget line

"56 = 1.8y+y \\


56 = 2.8y \\


y = \frac{56}{2.8} = 20 \\


x= 36 \\


U(x,y)_{max} = 4 \times 36^2 + 3 \times 36 \times 20 + 6 \times 20^2 \\


= 5184 + 2160 + 2400 = 9744"